News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 9 May 2003 8.00am EDT

The observing schedule was interrupted at 4:08pm on 8 May due to a high radiation event that activated the science instrument safing sequence (SCS 107). The high radiation was associated with an active coronal stream as Chandra approached the radiation belts. The loads have been replanned to start at 9:39am on May 9. The observation of Sco X-1 was impacted by the event (total 14Ks science time). Future loads will include an increased radiation belt pad time to try and avoid a similar trip.

A real-time procedure was completed on May 6 to perform a manual unload of momentum using thrusters A-1 (100%), A-2 and A-4. The unload was performed at an attitude that ensured a cooler starting temperature than prior anomalous unloads. Thruster performance was nominal during the unload and indicates that the thermal environment is related to the prior anomalous performance.

Real time procedures were completed on May 6 to perform a truncated ACA dark current calibration in support of a DSN Network Simplification Project test using DSS-59, and on May 8 to lower the ACA CCD temperature from -10 to -11 deg C to mitigate the increasing number of warm pixels. The ACA CCD will be lowered in a series of 1 deg steps to -15 deg C to allow for performance assessment.

The IRU-1 gyro-1 current remained at ~150mA (its yellow limit) with continued noisy behavior.

The observing schedule for next week is shown below and includes an observation of Mkn 279 coordinated with HST and FUSE.

------------------------------------------
RCW103                 ACIS-I       May 4
SDSS J222845-075755    ACIS-S
DOAR 21                ACIS-S/HETG  May 5
Radiation Belts                     May 6
PDS 456                ACIS-S/HETG
Sco X-1                ACIS-S/HETG  May 8
Radiation Belts
MXB 1659-298           ACIS-S       May 9
G11.2-0.3              ACIS-S
MRK 279                HRC-S/LETG   May 10
HD110073               ACIS-I
NGC 1800               ACIS-S

------------------------------------------

All spacecraft subsystems, with the exception of the MUPS thruster A-1 under investigation, continued to operate nominally.

Return to Status Reports